# Multi-objective Combinatorial Methodology for Nuclear Reactor Site Assessment: A Case Study for the United States

### v1

@@ 4.1 Reactor Siting Analysis in the United States
Figure 9: Best performing brownfield and coal sites of the joint dataset based on the non-dominated multi-objective combinatorial search.
The map includes 12 brownfield sites and 8 CPP sites, where each location is based on optimal combinations of the 22 objectives in the joint dataset. The best sites in this comparative study are located in Almeda County in CA, Yolo County in CA, Milwaukee County in WI, Plumas County in CA, Person County in NC. Analysis of objective contributions to the siting metric (score) reveals that each site excels in different dominant objectives impacting its performance. For instance, the model choses Almeda County, California site for its high electricity price, proximity to a significant number of nuclear R&D facilities, and optimal values across other key objectives. Compared to this, the Milwaukee County, Wisconsin site demonstrates stronger safety-related objectives, lower proximity advantages and distant to key objectives, and equivalent economic metrics. This example highlights clear trade-offs between the sites.
The method implemented in this study employs non-dominated sorting to identify the optimal locations. This approach selects locations that are not dominated by others based on the given criteria. The results of the method depend on the dataset used. In the brownfield study, the best site identified was in St. Lucie County, Florida. The objective contributions indicate that this site was selected due to the absence of state nuclear restrictions, the presence of a traditionally regulated energy market, the lack of protected lands, fault lines, landslide hazards, and open water, a peak ground acceleration lower than 0.3g, a slope less than 12 percent, proximity to a substation, and the availability of streamflow.
If a site were to offer at least one superior objective while maintaining the same favorable attributes described above, the St. Lucie County site would be dominated in most combinations. Since the introduction of CPP sites has eliminated St. Lucie County, Florida, as a viable option, we can infer that a CPP site dominates the St. Lucie County site.
The detailed siting metrics and objective values for the top CPP and brownfield sites are provided in Table 8 in the Appendix. The objective importance values of the best brownfield and coal sites in the joint dataset is given in Figure 10.
Figure 10: Summed normalized objective contributions ($\overrightarrow{SC}$) of the best-performing locations in the brownfield and coal site joint dataset. Further details about each site are given in Table 8.
@@
Upon examining the results of the joint brownfield and coal sites, we identify the top two performing sites as brownfield locations. In the top 20 sites, the number of brownfields and coal power plant (CPP) locations are close to each other, meaning that both site options are competitive. Notably, the CPP sites predominantly meet safety objectives, as they are generally located away from external hazards. Coal sites are also strategically positioned near electricity substations, streamflow resources, and major roads, providing an advantage in grid connectivity and infrastructure accessibility over brownfield sites. However, the brownfield sites span locations across all contiguous U.S. states. Given the higher number of brownfield sites relative to coal power plants, it is more likely that optimal locations with strong overall objectives can be identified among brownfields instead of CPPs.
@@
Our research methodology employs non-dominated sorting to identify locations that dominate across multiple objectives, without considering the specific magnitude of each objective. This approach prioritizes a location’s relative performance on individual objectives rather than their absolute values. Consequently, even if an objective holds the extreme (best or worst) value in the dataset, its specific magnitude is not factored into the sorting process. For instance, the highest-ranked location has several nearby hazardous facilities, yet it outperforms others due to advantageous factors such as state electricity imports, proximity to retiring facilities, nearby nuclear R&D centers, and distances to both electricity substations and transportation systems. This outcome underscores that certain important objectives in top locations are outweighed by the sheer number of other objectives. This illustrates how trade-offs become inherent to multi-objective optimization.
@@ 4.2 ConcNN Model Results
Based on the method presented in Section 3.2 and after collecting an enormous amount of information from our combinatorial search, we built a data-driven model to predict the NPP siting objectives, siting metric, and objective importance values based on the site location.
The model evaluations in this section use Mean Squared Error (MSE), Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and $R^{2}$ (coefficient of determination). MSE measures the average squared difference between predicted and actual values, penalizing larger errors more heavily. RMSE, the square root of MSE, expresses the error in the same units as the target variable, making it more interpretable. MAE calculates the average absolute differences, providing a more direct measure of typical error without emphasizing outliers. $R^{2}$ evaluates the proportion of variance in the target variable explained by the model, with values closer to 1 indicating better fit. We report these metrics in tables for the evaluated models.
The best ConcNN model after tuning consists of six layers of 1000 neurons in its first part, and two layers of 400 neurons in its second part. We set the learning rate to 1e-3. We select the test data proportion to be 20% of the total dataset (6,842 site locations). The test set is used to assess model performance through site locations not used during training. We set the number of epochs to 1,000 and the batch size to 128. The siting objective prediction ($Y1$) and siting metric and objective importance prediction ($Y2$) training metrics for the ConcNN model are given in Table 5. The model training loss is given in Figure 12. The best ConcNN architecture is sketched in Figure 11.
@@
Y1 Training 5.49058 2.34324 0.55280 0.85176
Y1 Test 7.6995 2.77481 0.58179 0.79706
Y2 Training 0.00018 0.01355 0.00657 0.76858
Y2 Test 0.00023 0.01522 0.00749 0.70403
Figure 12: The change of $Y_{L}$, $Y_{B}$, $Y_{2}$ training losses and the total training loss with epoch.
The metric results in Table 5 indicate that while the ConcNN model performs reasonably well, it lacks high accuracy in predicting site characteristics, with an $R^{2}$ on the test set ranging from 0.7 to 0.8 for its two output layers ($Y1$ and $Y2$), compared to the ideal $R^{2}$ of 1.0. This suggests that, despite being trained on a dataset of approximately 34,000 points, the problem remains challenging to model purely by machine learning. The primary difficulty arises because the model receives only four input features related to the site’s location and is tasked with predicting 22 site objectives. As a result, predictions may be particularly error-prone near state borders where site objectives objectives may exhibit significant variability in response to minor changes in input parameters. Furthermore, the lower accuracy in predicting $Y1$ propagates to $Y2$, leading to even lower $R^{2}$ values for the second output. Despite these limitations, the authors believe that the model can still offer rapid predictions of site objectives, site scores, and objective importance values that capture general trends across U.S. sites. However, its errors may be significant enough to limit its utility for precise site ranking.
@@ 4.3 LUT-NN Model Results
We present an assessment of LUT-NN, which replaces the first part of ConcNN with a lookup table and linear interpolation, in this section. The hyperparameter tuning for the LUT-NN was conducted over a broad range of values: the number of layers varied from 1 to 10, the number of neurons per layer ranged from 25 to 1,000, and the learning rate was tested between 1e-3 and 1e-5. The optimal model configuration identified consisted of six layers with neuron counts of [750, 750, 750, 700, 700, 650] in each layer, and an initial learning rate of 2e-4. We applied a learning rate decay factor of 0.92 with early stopping patience set to 25 epochs and trained the model for 1,000 epochs with a batch size of 32. To address overfitting, we applied L2 regularization with values ranging from 1e-2 to 1e-5, as well as dropout rates between 0.05 and 0.5, to each layer. Unfortunately, these strategies were not effective in improving prediction accuracy for locations that are absent from the lookup table.
The prediction metrics of LUT-NN for the $Y1$ values are shown in Table 6 for the test sites that exist in the dataset. As expected, the model perfectly predicts test sites. This is represented by the metrics in Table 6. However, to evaluate the $Y1$ metrics with LUT-NN for the site locations that do not exist in the lookup table, 15% of the existing lookup table is selected as the test data (which is similar to the concept used for testing ConcNN). By using the separated test data, we found the performance of the siting objective interpolation to be comparable to the predictions made by the first part of the ConcNN model (Y1). The test $R^{2}$ for the ConcNN is about 0.8 while it is 0.81 for the LUT-NN. This implies that the first part of the ConcNN could not learn more than what linear interpolation between the sites could learn.
We present the prediction metrics for LUT-NN on the $Y1$ values (22 site objectives) in Table 6 for test sites that already “exist” in the dataset. As expected, the model performs perfectly for these sites, reflected by the metrics in Table 6, as the model only needs to match the input data with the right site index. However, to assess the $Y1$ metrics for site locations not present in the lookup table, 15% of the existing lookup table was designated as test data, following a similar approach used to evaluate ConcNN. Using this separated test data, as in Table 6, the performance of LUT-NN in interpolating siting objectives was found to be comparable to the predictions made by the first part of the ConcNN model ($Y1$). Specifically, the test $R^{2}$ is approximately 0.8 for ConcNN and 0.81 for LUT-NN. This result suggests that the first stage of the ConcNN model does not capture more information than what can be achieved through linear interpolation between sites.
@@
$Y_{1}$ Test Set (Non-existing Site) 6.95979 2.63814 0.55284 0.81520
The most notable difference between the two models emerges when the site objectives ($Y1$) are used to predict the second set of outputs ($Y2$: site score and objective importance values). As noted earlier in Table 5, the $Y2$ metrics for ConcNN are even lower than those for $Y1$ with 20% test sites withheld from the full dataset. However, as shown in Table 7, when LUT-NN is used to predict $Y2$ for the same test sites as ConcNN, we see significant improvement. The training $R^{2}$ increases to nearly 0.99, while the test $R^{2}$ improves to approximately 0.88, compared to 0.7 for ConcNN. This demonstrates that using LUT-NN to predict site scores and objective importance values is considerably more accurate than relying solely on a neural network-based approach. As previously mentioned, we applied various regularization techniques to mitigate overfitting, and the results in Table 7 represent the best outcomes achieved.
Table 7: Lookup Table and Interpolation Informed Model Training and Test Metrics
@@
Y2 Training 6.314e-06 0.0025 0.0007 0.9912
Y2 Test 7.840e-05 0.0089 0.0028 0.8753
For both ConcNN and LUT-NN, the user is required only to input the location coordinates along with the FIPS codes for the county and state, which are available from the Federal Communications Processing System [^2] https://transition.fcc.gov/oet/info/maps/census/fips/fips.txt. The model then predicts NPP siting objectives, siting metrics, and objective importances by interpolating between more than 34,000 site locations available to the model. We designed these models to facilitate accessible and efficient retrieval of NPP siting information across locations in the contiguous U.S.. Based on our analysis,NN can be used for general site assessment without ranking, but the LUT-NN is much more reliable when predicting both site objectives for any site in the U.S. LUT-NN should be preferred for site ranking, as the site scores and objective importance values are more accurate than those predicted by ConcNN.
@@ 5 Concluding Remarks
This study introduces a novel, multi-objective combinatorial methodology for NPP site assessment and ranking that eliminates the need for analyst-defined weights, reducing potential bias. The study marks the first comprehensive evaluation of a vast number of potential nuclear reactor sites in the United States, considering both brownfield and coal sites, using a unified, flexible methodology. The proposed approach is adaptable and can be applied to other countries beyond the United States. Furthermore, we developed a machine learning model using the extensive dataset generated through this combinatorial process that enables rapid assessments of nuclear reactor site suitability across the U.S. This model requires only basic site location information (site coordinates, county, and state) from the user.
@@
We generated a comprehensive dataset encompassing raw siting objectives for NPPs in the United States for both brownfield and coal sites. Researchers can use this dataset in its entirety (22 objectives) or can focus on specific objectives relevant to their studies. It includes a wide range of socioeconomic, safety, and proximity factors. Notably, most of these siting characteristics are uncorrelated, preserving distinct information about various locations across the U.S.. We anticipate the brownfield NPP siting dataset to greatly enhance siting analysis and facilitate comparisons of site performance in subsequent research endeavors to aid in carbon-free transition.
@@
In this analysis, we identified several brownfield sites that present viable opportunities for the siting of NPPs. The findings, detailed previously, demonstrate that these sites possess the necessary socioeconomic, geographic, environmental, and proximity characteristics, and have advantages in terms of existing infrastructure and minimized land-use conflicts. Future research should continue to explore the implications of NPP siting on brownfield sites, specifically considering other factors, such as community acceptance to ensure holistic decision-making. To ensure the thorough evaluation and justification of the final siting decisions for the proposed NPPs, we expect that a detailed socio-techno-economic analysis is needed for the top-ranked sites found in this study.
@@
Our analysis also reveals a complex situation in the comparison of nuclear energy generation performance between CPPs and brownfield sites for NPP siting. While some CPPs demonstrate superior performance compared to brownfields, it is noteworthy that certain brownfield sites outperform CPPs. This variability underscores the complexity of energy site assessment. Given these findings, it is clear that the brownfields can compete with CPPs when socioeconomic, safety, and proximity characteristics are adequately considered. The findings of this research indicate that brownfields should be considered alongside coal sites for nuclear projects. Still, this research found that coal power plant sites outperformed most of the brownfield sites, which indicates that coal sites are a competitive option. This conclusion agrees with the previous research conducted in this area [19, 10] using methods distinct from those in this study.
@@
The dataset and models created in this study will be made available on a public Github repository following the completion of the peer review process. The authors will share the repository with the reviewers during an advanced stage of the review process.

### v2

@@ 4.1 Reactor Siting Analysis in the United States
Figure 9: Best performing coal and Brownfield sites of the joint dataset based on the non-dominated multi-objective combinatorial search.
@@
The map identifies six Brownfield sites and 14 CPP sites, each selected based on optimal combinations of 22 objectives from the joint dataset. Upon examining the results of the Brownfield and coal sites, we identify the top four performing sites as CPP locations. While CPPs represent a smaller subset of the overall dataset, their strong siting characteristics contribute to their superior performance. In the top 20 sites, the number of CPP locations are higher than Brownfield locations, meaning that CPP sites are more competitive. Notably, the CPP sites predominantly meet safety objectives, as they are generally located away from external hazards. Coal sites are also strategically positioned near electricity substations, streamflow resources, and major roads, providing an advantage in grid connectivity and infrastructure accessibility over Brownfield sites. Therefore, the prevalence of CPPs among the top 20 selected sites is both expected and justified. While they are not strategically positioned, the Brownfield sites span across all contiguous U.S. states. Given the higher number of Brownfield sites relative to coal power plants, it is expected that some locations comparable to the CPP sites can be identified among Brownfields.
The method implemented in this study selects locations that are not dominated by others based on the given criteria. Analysis of objective contributions to the siting score reveals that each site excels in different dominant objectives impacting its performance. The detailed siting score and objective values for the top CPP and Brownfield sites are provided in Table 8 in the Appendix. The objective importance values of the best Brownfield and coal sites in the joint dataset are given in Figure 10.
Figure 10: Summed normalized objective contributions ($\widehat{SC}$) of the best-performing locations in the Brownfield and coal site joint dataset. Further details about each site are given in Table 8.
@@
Our research methodology employs non-dominated sorting to identify locations that dominate across multiple objectives, without considering the specific magnitude of each objective. This approach prioritizes a location’s relative performance on individual objectives rather than their absolute values. Consequently, even if an objective holds the extreme (best or worst) value in the dataset, its specific magnitude is not factored into the sorting process. For instance, the highest-ranked location has a low electricity price, yet it outperforms others due to advantageous factors such as state electricity imports, safety objectives, proximity to retiring facilities, and distances to electricity substations and population centers. This outcome underscores that certain important objectives in top locations are outweighed by the sheer number of other objectives. This illustrates how trade-offs become inherent to multi-objective optimization.
The proposed methodology may unintentionally overlook crucial domain expertise that could better prioritize objectives based on practical constraints, such as safety or cost. For example, if all socioeconomic factors are positive and dominate the dataset, but a single safety objective is extremely low, the methodology might assign a high siting score to this location. However, an expert would quickly recognize that this site is unsuitable for nuclear power plant (NPP) siting. A similar issue arises with the current use of the “weighted sum method,” where even if the critical safety objective is assigned a higher-than-average weight, it may still be overshadowed by other objectives with lower-than-average weights, but whose collective impact is still significant. Thus, it is important to note that domain expertise is necessary to validate the results before final site selection, in both the proposed combinatorial method and the weighted sum method. Incorporating domain-specific constraints, hierarchical grouping of objectives, or expert post-hoc validation could help mitigate the risk of an outlier site being ranked too highly.
@@ 4.2 ConcNN Model Results
Based on the method presented in Section 3.2 and after collecting an enormous amount of information from our combinatorial search, we built a data-driven model to predict the NPP siting objectives, siting score, and objective importance values based on the site location.
Model evaluation in this section employs various regression metrics including Mean Squared Error (MSE), Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and $R^{2}$ (coefficient of determination). MSE measures the average squared difference between predicted and actual values, penalizing larger errors more heavily. RMSE, the square root of MSE, expresses the error in the same units as the target variable, making it more interpretable. MAE calculates the average absolute differences, providing a more direct measure of typical error without emphasizing outliers. $R^{2}$ evaluates the proportion of variance in the target variable explained by the model, with values closer to 1 indicating better fit. Mathematically, these four metrics are defined as follows:
MSE $=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$ RMSE $=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$ (10)
MAE $=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$ $\quad R^{2}$ $=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (11)
where $n$ is the number of samples in the test set (e.g., sites not used in training the model), $\hat{y}_{i}$ is the model prediction of sample $i$ (e.g., site score, objective values, objective importance values), $y_{i}$ is the true values for sample $i$ that correspond to $\hat{y}_{i}$ predictions, and $\bar{y}$ is the mean of such true values.
The best ConcNN model after tuning consists of seven layers of 600 neurons in its first part, and five layers of 950 neurons in its second part. We set the learning rate to 1e-3. We select the test data proportion to be 20% of the total dataset (6,895 site locations). The test set is used to assess model performance through site locations not used during training. We set the number of epochs to 2,000 and the batch size to 256. The siting objective prediction ($Y1$) and siting score and objective importance prediction ($Y2$) training metrics for the ConcNN model are given in Table 5. The model training loss is given in Figure 12. The best ConcNN architecture is sketched in Figure 11.
@@
Y1 Training 66485 257.848 44.8439 0.86388
Y1 Test 101831 319.110 51.3142 0.77205
Y2 Training 0.00028 0.00911 0.00313 0.92680
Y2 Test 0.00027 0.01654 0.00662 0.75634
Figure 12: The change of $Y_{L}$, $Y_{B}$, $Y_{2}$ training losses and the total training loss with the number of epochs for the ConcNN model.
The metric results in Table 5 indicate that while the ConcNN model performs reasonably well, it lacks high accuracy in predicting site characteristics, with an $R^{2}$ on the test set ranging from 0.75 to 0.75 for its two output layers ($Y1$ and $Y2$), compared to the ideal $R^{2}$ of 1.0. This suggests that, despite being trained on a dataset of approximately 34,000 points, the problem remains challenging to model purely by machine learning. The primary difficulty arises because the model receives only 4 input features related to the site’s location and is tasked with predicting 22 site objectives. As a result, predictions may be particularly error-prone near state borders where site objectives may exhibit significant variability in response to minor changes in input parameters. Furthermore, the lower accuracy in predicting $Y1$ propagates to $Y2$, leading to even lower $R^{2}$ values for the second output. Despite these limitations, the authors believe that the model can still offer rapid predictions of site objectives, site scores, and objective importance values that capture general trends across U.S. sites. However, its errors may be significant enough to limit its utility for precise site ranking.
@@ 4.3 LUT-NN Model Results
In this section, we present an assessment of LUT-NN, which replaces the first part of ConcNN with a lookup table and linear interpolation. The hyperparameter tuning for the LUT-NN was conducted over a broad range of values: the number of layers varied from 1 to 10, the number of neurons per layer ranged from 25 to 1,000, and the learning rate was tested between 1e-3 and 1e-5. During the hyperparameter tuning, a parameter grid has been generated for the defined parameter range. The best model was selected after the grid search was completed over all possible combinations of the parameters. After hyperparameter tuning, the optimal model configuration identified consisted of five layers with 950 neurons in each layer, and an initial learning rate of 2e-4. We applied a learning rate decay factor of 0.92 with early stopping patience set to 25 epochs and trained the model for 1,000 epochs with a batch size of 16. To address overfitting, we applied L2 regularization with values ranging from 1e-2 to 1e-5, as well as dropout rates between 0.05 and 0.5, to each layer. Unfortunately, these strategies were not effective in improving prediction accuracy for locations that are absent from the lookup table.
We present the prediction metrics for LUT-NN on the $Y1$ values (22 site objectives) in Table 6 for test sites that already “exist” in the dataset. As expected, the model performs perfectly for these sites, reflected by the metrics in Table 6, as the model only needs to match the input data with the right site index. However, the interpolation results heavily depend on the locations used in the lookup table. In order to assess the $Y1$ metrics for site locations not present in the lookup table, a 100-fold test set has been generated with 1% of the existing lookup table as test data and the remaining as the interpolation data. Using this separated test data, as shown in Table 6, the averaged performance of LUT-NN in interpolating siting objectives was found to be comparable to the predictions made by the first part of the ConcNN model ($Y1$). Specifically, the test $R^{2}$ is approximately 0.77 for the first part of ConcNN and 0.80 for LUT-NN. This result suggests that the first stage of the ConcNN model does not capture more information than what can be achieved through linear interpolation between sites.
@@
Objective ($Y_{1}$) Interpolation (Non-existing Site) 17.52051 4.06095 0.69053 0.80116
The test metrics (MSE, RMSE, and MAE) for interpolation in Table 6 differ significantly from those of the first-layer ConcNN (Y1). The interpolation table retrieves existing state information when site data is requested. Since this lookup table contains data from all contiguous U.S. states, it perfectly predicts state-dependent attributes. These attributes include key socioeconomic characteristics: (1) the number of nuclear restrictions in the state, (2) state electricity price, (3) state net electricity imports, (4) state nuclear-inclusive policy, (6) traditional regulation in the energy market, and (7) the 5-year average labor rate. Among these, (3) state net electricity imports and (7) the 5-year average labor rate have values that are considerably larger in magnitude compared to the other characteristics. Because the first-layer ConcNN exhibits errors in these columns while the lookup table does not, the MSE, RMSE, and MAE metrics between these methods differ significantly.
The most notable difference between the two models emerges when the site objectives ($Y1$) are used to predict the second set of outputs ($Y2$: site score and objective importance values). As noted earlier in Table 5, the $Y2$ metrics for ConcNN are even lower than those for $Y1$ with 20% test sites withheld from the full dataset. However, as shown in Table 7, when LUT-NN is used to predict $Y2$ for the same test sites as ConcNN, we see significant improvement. The training $R^{2}$ increases to approximately 0.97, while the test $R^{2}$ improves to 0.84 for the locations close to the lookup table sites and 0.79 for the locations far to the lookup table sites, compared to 0.75 for ConcNN. This demonstrates that using LUT-NN to predict site scores and objective importance values is considerably more accurate than relying solely on a neural network-based approach. As previously mentioned, we applied various regularization techniques to mitigate overfitting, and the results in Table 7 represent the best outcomes achieved.
Table 7: LUT-NN Training and Test Metrics when predicting $Y_{2}$ based on interpolated $Y_{1}$ values
@@
$Y_{2}$ Training 0.000027 0.005201 0.000878 0.973305
$Y_{2}$ Test (Without Y1 Interpolation Error) 0.000162 0.012728 0.004583 0.840029
$Y_{2}$ Test (With Y1 Interpolation Error) 0.000237 0.015380 0.004566 0.79412
For both ConcNN and LUT-NN, the user is required only to input the location coordinates along with the FIPS codes for the county and state, which are available from the Federal Communications Processing System [^2] https://transition.fcc.gov/oet/info/maps/census/fips/fips.txt. The model then predicts NPP siting objectives, siting metrics, and objective importances by interpolating between more than 34,000 site locations available to the model. We designed these models to facilitate accessible and efficient retrieval of NPP siting information across locations in the contiguous United States. Based on our analysis, ConcNN can be used for general site assessment without ranking, but the LUT-NN is much more reliable when predicting both site objectives for any site in the United States. LUT-NN should be preferred for site ranking, as the site scores and objective importance values are more accurate than those predicted by ConcNN.
One key advantage of these data-driven models, along with their rapid prediction capabilities, is their ability to support sensitivity analysis and uncertainty quantification—both of which require repeated model executions. These tasks are impractical with combinatorial search methods, as the process must be repeated whenever an objective is removed. However, with the LUT-NN model, users can exclude an objective (e.g., electricity substation distance) and assess its impact on the site score using techniques like one-at-a-time sensitivity analysis. Alternatively, they can conduct more comprehensive variance-based sensitivity analyses, such as those using Sobol indices, while keeping computational costs manageable. To maintain the focus of this work, we defer in-depth sensitivity analysis and uncertainty quantification to future studies.
The development of this machine learning-based NPP site selection algorithm has profound implications for the future of nuclear energy and infrastructure planning. By using data-driven decision-making built on data generated by unbiased combinatorial search, we minimize human bias, accelerate site selection, and enhance the overall process. The development of this method may help in receiving regulatory approvals, optimizing the costs of site selection, and ensuring that nuclear sites are strategically located for grid stability and resource efficiency. Additionally, it fosters public trust by promoting transparency in the selection process. On a global scale, this approach can aid both developed and emerging nations in efficiently identifying optimal sites for NPPs. Ultimately, this research paves the way for a more comprehensive evaluation of nuclear energy infrastructure crucial for meeting future energy demands and climate goals.
@@ 5 Concluding Remarks
This study introduces a novel, multi-objective combinatorial methodology for NPP site assessment and ranking that eliminates the need for analyst-defined weights, reducing potential bias. The study marks the first comprehensive evaluation of a vast number of potential nuclear reactor sites in the United States, considering both Brownfield and coal sites, using a unified, flexible methodology. The proposed approach is adaptable and can be applied to other countries beyond the United States. Furthermore, we developed a machine learning model using the extensive dataset generated through this combinatorial process that enables rapid assessments of nuclear reactor site suitability across the United States. This model requires only basic site location information (site coordinates, county, and state) from the user.
@@
We generated a comprehensive dataset encompassing raw siting objectives for NPPs in the United States for both Brownfield and coal sites. Researchers can use this dataset in its entirety (22 objectives) or can focus on specific objectives relevant to their studies. It includes a wide range of socioeconomic, safety, and proximity factors. Notably, most of these siting characteristics are uncorrelated, preserving distinct information about various locations across the United States.
@@
In this analysis, we identified some specific Brownfield sites that present viable opportunities for the siting of NPPs even when compared with the CPP sites. The findings, detailed previously, demonstrate that these specific sites possess the necessary socioeconomic, geographic, environmental, and proximity characteristics, and have advantages in terms of existing infrastructure and minimized land-use conflicts. Future research should continue to explore the implications of NPP siting on Brownfield sites, specifically considering other factors, such as community acceptance to ensure holistic decision-making. To ensure the thorough evaluation and justification of the final siting decisions for the proposed NPPs, we expect that a detailed socio-techno-economic analysis is needed for the top-ranked sites found in this study including both the coal and Brownfield sites.
@@
Our analysis also reveals a clear situation in the comparison of nuclear energy generation performance between CPPs and Brownfield sites for NPP siting. While some Brownfields demonstrate comparable performance compared to CPPs, most CPP sites outperform the entire Brownfield dataset. This finding underscores the justification of the selected CPP sites when considered in the context of energy site assessment. This conclusion agrees with the previous research conducted in this area [22, 13] using methods distinct from those in this study. Given these findings, it is also shown that some exceptional Brownfields can compete with CPPs when socioeconomic, safety, and proximity characteristics are adequately considered. The findings of this research indicate that Brownfields should be considered alongside coal sites for nuclear projects.
One limitation observed in this work is that neural network models still have room for improvement, particularly in their ability to generalize to new sites. This challenge likely stems from data variability and model overfitting. The current dataset may be insufficient to ensure broad generalization across all states, highlighting the need for more data to enhance accuracy and robustness. Future research should explore advanced regularization techniques and improved data acquisition methods to address these issues. Additionally, further work on data-driven models should include sensitivity analysis and uncertainty quantification to assess their robustness against both minor and major perturbations in input data.
@@
The dataset and models created in this study are available on the public Github repository: https://github.com/aims-umich/NPP_Siting.